(1) The table below shows the calories, fat, and carbohydrates per ounce for three brands of cereal, and the total amount of calories, fat, and carbohydrates required for a special diet. Calories Fat Carbohydrates Cereal Brand X 50 0 22 Cereal Brand Y 108 0.1 25.5 Cereal Brand Z 127 5.5 18 Total Required 393 5.7 91 (i) Using the information in the table above, write a system of three equations. (ii) Write a system in matrix form Ax b ๏€ฝ . (ii) Using the Cramer's Rule or Inverse Method, find the amount of each brand of cereal that will give the level of nutrition required.ย 


1
Expert's answer
2021-11-11T08:53:04-0500

Using cramer's Rule;


"\begin{pmatrix}
 50&108& 127\\
 0&0.1 & 5.5\\
22&25.5&18
\end{pmatrix}\begin{pmatrix}
 x \\
 y \\
z
\end{pmatrix}=\begin{pmatrix}
 393 \\
 5.7\\
91
\end{pmatrix}"



Let "D=" "\begin{pmatrix}
 50&108 & 127 \\
 0 & 0.1&5.5\\
22&25.5&18
\end{pmatrix}"



"D_x=\begin{pmatrix}
 393& 108&127\\
 5.7&0.1 & 5.5\\
91&25.5&18
\end{pmatrix}"



"D_y=\begin{pmatrix}
 50&393 & 127 \\
 0 & 5.7&5.5\\
22&91&18
\end{pmatrix}"



"D_2=





\begin{pmatrix}
 50&108 & 393 \\
 0&0.1 & 5.7\\
22&25.5&91
\end{pmatrix}"


Det "D=50\begin{vmatrix}
 0.1 & 5.5 \\
 25.5 & 18
\end{vmatrix}-108\begin{vmatrix}
 0& 5.5 \\
 22 & 18
\end{vmatrix}+127\begin{vmatrix}
 0 & 0.1\\
 22 & 25.5
\end{vmatrix}"



"=50(-138.45)-108(-121)+127(-2.2)"

"=5866.1"



Det "D_x=393\begin{vmatrix}
 0.1& 5.5 \\
 25.5& 18
\end{vmatrix}-108\begin{vmatrix}
 5.7& 5.5 \\
 91& 18
\end{vmatrix}+127\begin{vmatrix}
 5.7 & 0.1\\
 91 & 25.5
\end{vmatrix}"


"=393(-138.45)-108(-397.9)+127(136.25)"

"=5866.1"



Det "D_y=50\begin{vmatrix}
 5.7 & 5.5 \\
 91& 18
\end{vmatrix}-393\begin{vmatrix}
 0& 5.5 \\
 22& 18
\end{vmatrix}+127\begin{vmatrix}
 0& 5.7 \\
 22 & 91
\end{vmatrix}"


"=50(-397.9)-393(-121)+127(-125.4)"

"=11732.2"



Det "D_2=50\begin{vmatrix}
 0.1 & 5.7 \\
 25.5 & 91
\end{vmatrix}-108\begin{vmatrix}
 0& 5.7 \\
 22 & 91
\end{vmatrix}+393\begin{vmatrix}
 0& 0.1\\
 22& 25.5
\end{vmatrix}"


"=50(-136.25)-108(-125.4)+393(-22)"

"=5866.1"




"x=\frac{|D_x|}{|D|}=\frac{5866.1}{5866.1}=1,"


"y=\frac{|D_y|}{|D|}=\frac{11732.2}{5866.1}=2"

"z=\frac{|D_z|}{|D|}=\frac{5866.1}{5866.1}=1"



Amount of each brand of cereal that will give the level of nutrition required


Cereal Brand "X=1" ounce

Cereal Brand "Y=2" ounces

Cereal Brand "Z=1" ounce